# Approximation of arbitrarily high-order PDEs by first-order hyperbolic relaxation

### v1

@@ Abstract
We present a framework for constructing a first-order hyperbolic system whose solution approximates that of a desired higher-order evolution equation. Constructions of this kind have received increasing interest in recent years, and are potentially useful as either analytical or computational tools for understanding the corresponding higher-order equation. We perform a systematic analysis of a family of linear model equations and show that for each member of this family there is a stable hyperbolic approximation whose solution converges to that of the model equation in a certain limit. We then show through several examples that this approach can be applied successfully to a very wide range of nonlinear PDEs of practical interest.
@@ 1 Introduction
Virtually any textbook on ordinary differential equations (ODEs) includes the observation that higher-order ODEs can be rewritten as systems of first-order ODEs. This can facilitate the analysis of ODE solutions by placing all ODEs in a unified framework, and it is particularly useful in the realm of numerical approximation, since it allows numerical analysts to focus exclusively on methods for first-order problems. High-order partial differential equations (PDEs) can also superficially be written in first-order form through the introduction of auxiliary variables. However, this is less useful since the theory and analysis of first-order PDEs is developed only for hyperbolic systems, but the first-order form of a higher-order PDE is in general not hyperbolic.
In this work we present a general technique for constructing a system of first-order hyperbolic PDEs (referred to as a hyperbolic relaxation, or hyperbolization) whose solution approximates that of a given system of higher-order evolution PDEs. We begin by presenting some motivating examples in Sections 1.2-1.3. We present these known examples in a way that suggests a more general approach but also raises some key questions. We then develop the general idea by studying a simple family of linear PDEs, in Section 2, where we answer the questions raised previously by determining the necessary and sufficient conditions for stability of the hyperbolization. This leads to a unique hyperbolization for each member of the family of high-order PDEs. In Section 3 we prove a convergence result relating the solution of the stable hyperbolic relaxation constructed in Section 2 and the original high-order PDE. In Section 4 we provide some examples suggesting that our approach can be successfully applied in a straightforward way to a wide range of nonlinear PDEs, including complex-valued equations, systems of equations, equations with mixed derivatives, and equations with up to (at least) 4th-order derivatives. We conclude in Section 5 with a discussion of important open questions.
@@ 1.1 Prior work and motivation
The idea of approximation by a hyperbolic system dates back at least to the works of Cattaneo and Vernotte [5, 25], who constructed a hyperbolic relaxation of the heat equation. Much more recently, Toro & Montecinos [24] gave a general approach to and analysis of hyperbolization of second-order advection-reaction-diffusion equations, including a detailed study of the efficiency of the numerical solution of the hyperbolized model compared to the original model. Mazaheri et. al. [21] provided a general approach to hyperbolization of scalar PDEs that include first-, second-, and third-order spatial derivatives (i.e., advection-diffusion-dispersion equations), and demonstrated it through application to the KdV and Burgers equations. Many other hyperbolized versions of PDE models have been proposed, e.g. for Korteweg-de Vries (KdV) [3], Benjamin-Bona-Mahoney (BBM) [15], Serre-Green-Naghdi (SGN) [13, 12, 2], multilayer dispersive shallow water equations [7], the compressible Navier-Stokes equations [20], the Euler-Korteweg equation (in hydrodynamic form) [10], and a range of additional dispersive wave equations [1, 16, 11, 9, 6, 14], as well as elliptic equations [22].
@@
The approach introduced by Jin and Xin [18] is also referred to as hyperbolic relaxation, but it differs fundamentally from the concept dicsussed herein. In Jin-Xin relaxation, the relaxation terms are usually algebraic, the starting point is a first-order nonlinear hyperbolic system, and the result is a dissipative approximation. Here we start from higher-order (not hyperbolic) PDEs, we use relaxation terms that involve differential operators, and the resulting approximation is non-dissipative if the original problem is non-dissipative. See Section 1.2.1 for discussion of one connection between these ideas.
@@ 1.2 The Heat Equation
We see that $\omega_{+}(k)\approx-ik^{2}$, so that the dispersion relation of (1) is approximately preserved. The accuracy of the approximation is seen to improve as $\tau\to 0$. Notice that $\omega_{-}$ has negative real part proportional to $\tau^{-1}$, so there is a spurious mode but it decays quickly if $\tau$ is small. Note also that if $14k^{2}\tau>$ 1, then $\omega$ has a real part and a negative imaginary part. This means that solutions are stable but we should expect to see wave-like behavior for high wavenumbers or when $\tau$ is large. Note that if we take $\tau<0$, there exist exponentially growing solutions.
@@ 2.1.1 Low-wavenumber stability
For $k=0$, the dispersion relation (19) reduces to
@@ 2.1.2 High-wavenumber stability
As $|k|\to\infty$, the dispersion relation tends to the solution of
$\det(\omega I-kA)$ $=0,$ (22)
i.e. the dispersion relation of the homogeneous hyperbolic system $q_{t}+Aq_{x}$, whose solutions are stable if and only if $A$ has only real eigenvalues.
@@ 2.2 The unique stable hyperbolization
Having applied these necessary conditions, we can now ask if the system (17) with (26) is indeed stable for intermediate values of $k$ and for all values of $\tau$.
@@ 2.3 General Linear Scalar Evolution PDEs
2.4 Further generalizations
In Section 4 we will apply the approach above to more general PDEs, including nonlinear equations and equations with mixed space-time derivatives. For nonlinear problems, it is natural to require stability of the linearized dispersion relation, which leads again to the choice of signed permutation matrix determined above. For problems with mixed space- and time-derivatives, we will again use linear stability as our guiding principle. A detailed investigation of the hyperbolization of such PDEs, and of systems of PDEs, is left to future work.
@@ 3 Accuracy of the hyperbolic approximation
$\tau^{-1}>\frac{\partial_{x}^{4}u}{\partial_{x}^{2}u}$
in order to ensure that the error terms are small compared to the leading order terms. For the initial data shown in Figure 1, this suggests one should take $\tau^{-1}>60$, which is in general agreement with the results shown. As time advances, the ratio of derivatives becomes smaller, and one correspondingly observes improved agreement between the solutions (even those with large values of $\tau$) after long enough time.
@@
Consider the one-dimensional Cauchy problem for (12) with initial data
@@
Let $q$ denote the solution of the hyperbolization (27)-(28), and let $T$ be given such that $0\leq T<\infty$. Then
$\|q_{0}(x,T)-u(x,T)\|={\mathcal{O}}(\tau).$ (37)
Theorem 3 shows that the Fourier transform of the hyperbolized solution converges to that of the original problem pointwise. To prove convergence of the solution itself would require more detailed estimates of the difference between the two as a function of the wavenumber $k$. Such an extension, as well as extension to more general PDEs, is left to future work.
3.1 Proof of Theorem 3
@@
Property 5: This follows from the fact that $L_{\tau}$ is diagonalizable with distinct eigenvalues; see [19, Thm. 8, p. 130]. ∎
@@ 4 Examples
In the following examples, we use a Fourier pseudospectral collocation method in space and explicit 3rd-order Runge-Kutta integration in time. The code to reproduce each example is available online [^2] http://github.com/ketch/hyperbolization-RR.
@@ 4.2 The Camassa-Holm Equation
$\partial_{t}u-\partial_{t}\partial_{x}^{2}u+3u\partial_{x}u-2(\partial_{x})(u\partial_{x}^{2}u)+u\partial_{x}^{3}u=0,\ \text{on}\ [x_{L},x_{R}]\times(0,T]\;,$ (43)
with initial data $u(x,0)=\left(\frac{\pi}{2}\right)e^{x}-2\sinh(x)\arctan(e^{x})-1$ [4], and the periodic boundary condition $u(x_{L},t)=u(x_{R},t)$. Figure 7 in [4] illustrates that the solution of this problem. The initial parabolic pulse steepens and then breaks into a train of peakon solitons, which move at speeds proportional to their amplitudes. This example of hyperbolization is interesting in that it contains two different third-order derivative terms, one of which is nonlinear. Introducing equations for $q_{1}$ and $q_{2}$ in the same way as for the KdV equation, we write
@@
This system is not in the usual hyperbolic form, since the first equation contains time derivatives of both $q_{0}$ and $q_{2}$. This is easily remedied by adding $\tau^{-1}$ times the last equation to the first, leading to the hyperbolized CH (CHH) system
$\partial_{t}q_{0}+3q_{0}\partial_{x}q_{0}-2\partial_{x}q_{0}\partial_{x}q_{1}-q_{0}\partial_{x}q_{2}-\tau^{-1}(\partial_{x}q_{0}-q_{1})$ $=0$ (44a)
@@ 4.3 The Kuramoto-Sivashinsky Equation
with $\tau>0$. Solutions of (45) are known to exhibit both periodic behavior and chaotic behavior, depending on the problem setup. A preliminary study of solutions of (46) shows that it behaves similarly, yielding periodic or chaotic solutions under similar circumstances. The solutions of (46) are observed to converge to those of (45) as $\tau\to 0$, when the solution is not chaotic. Under conditions in which (45) becomes chaotic, that of (46) seems to do the same, and the time to onset of chaos is nearly the same.
@@ 5 Discussion
While the first work on hyperbolization of high-order PDEs dates back to the 1950s, work in this area has accelerated in the last 15 years, during which a variety of techniques in this vein have been applied to a number of different models. Some of these recent works can be seen as particular cases of the general technique discussed herein. Awareness and understanding of hyperbolization as a general tool is likely to facilitate its application to an increasing number of models.
Herein we have shown for the first time that a stable hyperbolization exists for PDEs of arbitrarily high order. We have focused on scalar evolution equations, in order to narrow the scope enough to facilitate development of a comprehensive theory. But a wide variety of other PDE models – including elliptic problems, evolution problems with hyperbolic constraints, and general systems of high-order PDEs – are susceptible to a similar treatment.
@@
Since hyperbolic PDEs admit similarity solutions (Riemann solutions) that can often be solved by semi-analytical means, the study of the Riemann problem for hyperbolized systems might yield novel insights. For one existing study of this kind, see [17].

### v2

@@ Abstract
We present a framework for constructing a first-order hyperbolic system whose solution approximates that of a desired higher-order evolution equation. Constructions of this kind have received increasing interest in recent years, and are potentially useful as analytical or computational tools for understanding the corresponding higher-order equation. We perform a systematic analysis of a family of linear model equations and show that for each member of this family there is a stable hyperbolic approximation whose solution converges to that of the model equation in a certain limit. We then show through several examples that this approach can be applied successfully to a very wide range of nonlinear PDEs of practical interest.
@@ 1 Introduction
Partial differential equations (PDEs) can be broadly grouped into important classes (e.g. hyperbolic, parabolic, elliptic) that share certain mathematical properties. Equations in different classes require different methods of analysis and numerical solution. In some cases it is known to be possible to approximate the solutions of a PDE in one class by solutions of a PDE in a different class. For instance, Cattaneo and Vernotte [5, 25], constructed a hyperbolic approximation (or relaxation) of the heat equation. While relatively little work was done on such approximations in subsequent decades, they have lately received a great deal of attention [1, 16, 24, 21, 13, 22, 19, 10, 12, 7, 2, 15, 3, 11, 9, 6, 14]. In each of the works just cited, the authors develop a hyperbolic approximation for a specific higher-order PDE, often with the goal of removing numerical stiffness. The techniques used in developing these equations appear to be specific to each individual equation. The spatial order of all the PDEs studied so far in this way is at most three.
In this work we present a general technique for constructing a system of first-order hyperbolic PDEs (referred to as a hyperbolic relaxation, or hyperbolization) whose solution approximates that of a given system of higher-order evolution PDEs. We begin by presenting some motivating examples in Sections 1.2-1.3. We present these known examples in a way that suggests a more general approach that unifies some of the existing work in this area and provides a hyperbolization for equations of arbitrary order. A key question that arises is the well-posedness of the Cauchy problem for the resulting hyperbolic system (i.e., the stability of the dispersion relation). In Section 2 we study a family of linear PDEs of arbitrary order, for which we derive the necessary and sufficient conditions for stability of the hyperbolization. This leads to a unique hyperbolization for any high-order evolution PDE, which is provably stable for linear PDEs in the family we study. In Section 3 we prove a convergence result relating the solution of the stable hyperbolic relaxation constructed in Section 2 and the original high-order PDE. We conjecture that the general hyperbolization we have constructed is also stable for other high-order linear and nonlinear PDEs (whenever the original equation is stable). In Section 4 we provide some examples supporting and extending this conjecture by applying our approach to various nonlinear PDEs, including complex-valued equations, systems of equations, equations with mixed derivatives, and equations with up to (at least) 4th-order derivatives. We conclude in Section 5 with a discussion of important open questions.
We emphasize that the main goal of the present work is not to provide further motivation for hyperbolization, but to dramatically increase its range of applicability and improve our understanding of it by tying together the numerous existing ad hoc examples through a unified general approach.
@@ 1.1 Prior work and motivation
Let us review in more detail the existing literature on hyperbolization and the proposed motivation for its use. Toro & Montecinos [24] gave a general approach to and analysis of hyperbolization of second-order advection-reaction-diffusion equations, including a detailed study of the efficiency of the numerical solution of the hyperbolized model compared to the original model. Mazaheri et. al. [21] provided a general approach to hyperbolization of scalar PDEs that include first-, second-, and third-order spatial derivatives (i.e., advection-diffusion-dispersion equations), and demonstrated it through application to the KdV and Burgers equations. Many other hyperbolized versions of PDE models have been proposed, e.g. for Korteweg-de Vries (KdV) [3], Benjamin-Bona-Mahoney (BBM) [15], Serre-Green-Naghdi (SGN) [13, 12, 2], multilayer dispersive shallow water equations [7], the compressible Navier-Stokes equations [19], the Euler-Korteweg equation (in hydrodynamic form) [10], and a range of additional dispersive wave equations [1, 16, 11, 9, 6, 14], as well as elliptic equations [22].
@@
The approach introduced by Jin and Xin [17, 20] is also referred to as hyperbolic relaxation, but it differs fundamentally from the concept dicsussed herein. In Jin-Xin relaxation, the relaxation terms are usually algebraic, the starting point is a first-order nonlinear hyperbolic system, and the result is a dissipative approximation. Here we start from higher-order (not hyperbolic) PDEs, we use relaxation terms that involve differential operators, and the resulting approximation is non-dissipative if the original problem is non-dissipative. See Section 1.2.1 for discussion of one connection between these ideas.
@@ 1.2 The Heat Equation
We see that $\omega_{+}(k)\approx-ik^{2}$, so that the dispersion relation of (1) is approximately preserved. The accuracy of the approximation is seen to improve as $\tau\to 0$. Notice that $\omega_{-}$ has negative real part proportional to $\tau^{-1}$, so there is a spurious mode but it decays quickly if $\tau$ is small. Note also that if $14k^{2}\tau>$ 1, then $\omega$ has a real part and a negative imaginary part. This means that solutions are stable but we should expect to see wave-like behavior for high wavenumbers or when $\tau$ is large.
@@ 2.1.1 Low-wavenumber stability
A necessary condition for stability at low wavenumbers is stability for $k=0$. In that case, the dispersion relation (19) reduces to
@@ 2.1.2 High-wavenumber stability
Dividing (19) by $k^{m}$ and taking the limit $k\to\infty$, we obtain the high-wavenumber limiting dispersion relation
$\det(c(k)I-A)$ $=0,$ (22)
where $c(k)=\omega(k)/k$ is the phase velocity. This is just the dispersion relation of the homogeneous hyperbolic system $q_{t}+Aq_{x}$, whose solutions are stable if and only if $A$ has only real eigenvalues.
@@ 2.2 The unique stable hyperbolization
These necessary conditions, which we obtained by considering the stability for both low and high wave numbers, turn out to be sufficient for the stability of the system (17) for all wave numbers $k$ and all $\tau>0$ The result is summarized in the following theorem.
@@
Note that in the limit $\tau=0$ , the relaxation system (17) simply reduces to the original PDE (12).
@@ 2.3 General Linear Scalar Evolution PDEs
The theory developed in this section is based on analysis of linear PDEs only. In Section 4 we will apply the approach above to more general PDEs, including nonlinear equations and equations with mixed space-time derivatives. For nonlinear problems, one expects that stability of the linearized problem is at least a necessary condition; this leads again to the choice of signed permutation matrix determined above. For problems with mixed space- and time-derivatives, we will again use linear stability as our guiding principle. A detailed investigation of the hyperbolization of such PDEs, and of systems of PDEs, is left to future work.
@@ 3 Accuracy of the hyperbolic approximation
$\tau^{-1}>\frac{\max_{x,t}\left|\partial_{x}^{4}u\right|}{\max_{x,t}\left|\partial_{x}^{2}u\right|}$
in order to ensure that the error terms are small compared to the leading order terms. For the initial data shown in Figure 1, this suggests one should take approximately $\tau^{-1}>60$, which is in general agreement with the results shown. As time advances, the ratio of derivatives becomes smaller, and one correspondingly observes improved agreement between the solutions (even those with large values of $\tau$) after long enough time.
In the next theorem we reference the following space of functions:
$\mathcal{F}_{n}$ =$=\left\{v(x)=\sum_{j=1}^{n}\hat{v}_{j}e^{ik_{j}x}:\hat{v},k_{j}\in\mathbb{R}\right\}$ $n<\infty.$ (36)
@@
Consider the one-dimensional Cauchy problem for (12) with initial data $u(x,t=0)\in\mathcal{F}_{n}$ for some $n$. Let $q$ denote the solution of the hyperbolization (27)-(28), and let $T$ be given such that $0\leq T<\infty$. Then
$\|q_{0}(x,T)-u(x,T)\|_{\infty}={\mathcal{O}}(\tau).$ (37)
Theorem 3 shows that the Fourier transform of the hyperbolized solution converges to that of the original problem pointwise. To prove convergence of the solution itself (or for a broader class of functions) would require more detailed estimates of the difference between the two as a function of the wavenumber $k$. Such an extension, as well as extension to more general PDEs, is left to future work.
3.1 Proof of Theorem 3
Due to linearity, it is sufficient to prove this result for the initial data
@@ 3.1 Proof of Theorem 3
Property 5: This follows from the fact that $L_{\tau}$ is diagonalizable with distinct eigenvalues; see [18, Thm. 8, p. 130]. ∎
@@ 4 Additional Nonlinear Examples
In the numerical solutions for these examples, we use a Fourier pseudospectral collocation method in space and explicit 3rd-order Runge-Kutta integration in time. The code to reproduce each example is available online [^2] http://github.com/ketch/hyperbolization-RR.
@@ 4.2 The Camassa-Holm Equation
$\partial_{t}u-\partial_{t}\partial_{x}^{2}u+3u\partial_{x}u-2\partial_{x}u\partial_{x}^{2}u-u\partial_{x}^{3}u=0,\ \text{on}\ [x_{L},x_{R}]\times(0,T]\;,$ (44)
with initial data $u(x,0)=\left(\frac{\pi}{2}\right)e^{x}-2\sinh(x)\arctan(e^{x})-1$ [4], and the periodic boundary condition $u(x_{L},t)=u(x_{R},t)$. Figure 7 in [4] illustrates the solution to this problem. The initial parabolic pulse steepens and then breaks into a train of peakon solitons, which move at speeds proportional to their amplitudes. This example of hyperbolization is interesting in that it contains two different third-order derivative terms, one of which is nonlinear. Introducing equations for $q_{1}$ and $q_{2}$ in the same way as for the KdV equation, we write
@@
This system is not in the usual hyperbolic form, since the first equation contains time derivatives of both $q_{0}$ and $q_{2}$, and the term $2\partial_{x}q_{0}\partial_{x}q_{1}$, which is a product of two first-order partial derivatives. The time derivative $\partial_{t}q_{2}$ can be eliminated by adding $\tau^{-1}$ times the last equation to the first equation. Additionally, the term $2\partial_{x}q_{0}\partial_{x}q_{1}$ can be replaced by either $2q_{1}\partial_{x}q_{1}$ or $2q_{2}\partial_{x}q_{0}$. We chose the latter, which leads to the hyperbolized CH (CHH) system
$\partial_{t}q_{0}+3q_{0}\partial_{x}q_{0}-2q_{2}\partial_{x}q_{0}-q_{0}\partial_{x}q_{2}-\tau^{-1}(\partial_{x}q_{0}-q_{1})$ $=0$ (45a)
@@ 4.3 The Kuramoto-Sivashinsky Equation
with $\tau>0$. Solutions of (46) are known to exhibit both periodic behavior and chaotic behavior, depending on the problem setup. A preliminary study of solutions of (47) shows that it behaves similarly, yielding periodic or chaotic solutions under similar circumstances. The solutions of (47) are observed to converge to those of (46) as $\tau\to 0$, when the solution is not chaotic.
@@ 5 Discussion
PDEs are perhaps the most widely used form of mathematical model, and herein we have shown that a very wide range of PDEs can be approximately transformed into a completely different class of PDE, potentially enabling their solution and analysis by different techniques than what have so far been applied to them.
While the first work on hyperbolization of high-order PDEs dates back to the 1950s, development in this area has accelerated in the last 15 years, during which a variety of techniques in this vein have been applied to a number of different models. Some of these recent works can be seen as particular cases of the general technique discussed herein. Awareness and understanding of hyperbolization as a general tool is likely to facilitate its application to an increasing number of models.
Herein we have shown for the first time that a stable hyperbolization exists for PDEs of arbitrarily high order. Our analysis has emphasize a particular family of linear PDEs, but the resulting approach provides a clear and systematic way to hyperbolize any scalar evolution equation (linear or nonlinear), and we have found this approach to be effective for every equation to which we have applied it. We have focused on scalar evolution equations, in order to narrow the scope enough to facilitate development of a comprehensive theory. But a wide variety of other PDE models – including elliptic problems, evolution problems with hyperbolic constraints, and general systems of high-order PDEs – are susceptible to a similar treatment.